I need an equation which models how old, in years, you will be when you are 3 times as old as the younger person. Ages: 42 and 15
Thanks for your help

Answer by ankor@dixie-net.com(22740)

I need an equation which models how old, in years, you will be when you are 3
times as old as the younger person. Ages: 42 and 15
:
Let x = no. of years ago when 42 yr old was 3 times 15 yr old
:
42 - x = 3(15 - x)
42 - x = 45 - 3x
-x + 3x = 45 - 42
2x = 3
x = 1.5 yrs ago
then
he will be 42 - 1.5 = 40.5 yrs old
and
child will be 15 - 1.5 = 13.5 yrs
:
Check: 3(13.5) = 40.5